HUBBLE SPACE TELESCOPE DAILY REPORT #2521 PERIOD COVERED: 0000Z (UTC) 12/15/99 - 0000Z (UTC) 12/16/99 Daily Status Report as of 350/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: Since HST remains in Zero Gyro Sun Point (ZGSP), no science observations were performed. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: No activity 2.2 FHST Updates: No activity 2.3 Operations Notes: The HST is in ZGSP. Beginning at noon, the regular flight operations team was joined by the numerous personnel that will be supporting SM3A. Using Operations Note #661, the time-based recharge ratio charge/recharge buffers were reset at 349/1516Z. The NSSC-1 status buffer was dumped at 349/2235Z per ROP NS-3. At 349/2356Z, the limits for the electronics and bearings for reaction wheels 1-4, and for the PSEA MCE , were set to values appropriate for SM3A. FGS-3 was put back into the default mode at 350/0216Z as directed by an operations request. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Servicing Mission 3A, currently still on schedule for a December 16, 1999 launch. The launch window opens at 351/0218Z. /CAW